there are three problems i need help
2+i=
3+2i+2+i=
5+2i.3+i=

Answer by Alan3354(31518)

2+i=
2 + 1
---------
3+2i+2+i=
5 + 3i
---------
5+2i.3+i=
Not sure what that means.